OpenStudy (anonymous):

What is the simplified form of the following expression ? 2x^2y+3x^2+4y+3x^2y+2y

jhonyy9 (jhonyy9):

so than you see you can assuming first with 4rd and 3rd with the last term can you make it ?

jhonyy9 (jhonyy9):

i will help you courage

jhonyy9 (jhonyy9):

ok like a start 2x^2y +3x^2y = 5x^2y do you understand it now ? continue please

OpenStudy (anonymous):

I do not understand what you mean by "so than you see you can assuming first with 4rd and 3rd with the last term can you make it ?"

jhonyy9 (jhonyy9):

so 4y +2y = ?

jhonyy9 (jhonyy9):

=(4+2)y = ?

OpenStudy (anonymous):

6Y ?

jhonyy9 (jhonyy9):

yes sure is right do you can continue it ?

jhonyy9 (jhonyy9):

so now there will be 5x^2y +3x^2 +6y yes ?

jhonyy9 (jhonyy9):

so now you can factorizing out the x^2(5y +3) +6y

OpenStudy (anonymous):

x^2(11y +3) i think

jhonyy9 (jhonyy9):

no because 5y is inside parentheses

OpenStudy (anonymous):

yeah but u cant add the 3 because it does not have a Y so dont u add the other one that has a y ?

jhonyy9 (jhonyy9):

yes what i have wrote above will be simplified form

OpenStudy (anonymous):

Oh that was the answer i thought it could be simplified more :/ Thank you so much for the help!

jhonyy9 (jhonyy9):

was my pleasure bye

OpenStudy (anonymous):

see ya and ty for the metal :)